OpenStudy (anonymous):

integrate (x^2)/(cos^2 x)

OpenStudy (abb0t):

\[\int\limits \frac{ x^2 }{ \cos^2(x) } dx\] is this correct?

OpenStudy (anonymous):

Yup.. should i factorise the square 1st? or using substitution?

OpenStudy (abb0t):

Hm, this one is tricky. Have you tried integrating by parts? Changing cos to sec?

OpenStudy (anonymous):

U mean integrate \[\int\limits_{}^{} x ^{2} \sec ^{2}x dx\] then let \[u = x ^{2}\] \[v = \int\limits_{}^{} \sec ^{2}x\] So,will hv \[x ^{2}\tan x - \int\limits_{}^{} 2x \tan x dx \]

OpenStudy (abb0t):

Yep. Use inegration by parts three times. It should work out.

OpenStudy (abb0t):

I will help you out by telling you for the second integration\[v = \ln|\cos(x)| \]

OpenStudy (abb0t):

I think it's best if you just memorize the integral of tan(x) because it's used often in integral calculus courses and physics courses.

OpenStudy (anonymous):

So \[\int\limits_{}^{}2x \tan x dx = -2x \ln \cos x +2 \int\limits_{}^{} \ln \cos x dx\] then,i have to integrate again for the lncosx,is it?

OpenStudy (abb0t):

Ah. It all worked out in my head, but now that I see it written down, it doesn't look right. Haha. Um... have you learned how to integrate using complex exponentials? Haha. I actually am stump'd now. Haha. sorry for the confusion.

OpenStudy (abb0t):

My second suggestion was going to be. \[\int\limits x^2(1-\tan^2(x))dx\]

OpenStudy (anonymous):

Haha. Its ok anyway. The question actually sounds like this \[\int\limits_{}^{} \frac{ x \sin 2x + x ^{2} }{ \cos ^{2}x } dx\] and i separate it into two parts. so \[\int\limits_{}^{} \frac{ x ^{2} }{ \cos ^{2} x}\] is the subpart of it.

OpenStudy (abb0t):

Oh, um, well, this is just my train of thought, not sure if it's right though.. \[\int\limits \frac{ x(2\sin(x)\cos(x))+x^2 }{ \cos^2(x) }dx \] Then you get: \[\int\limits \frac{ 2xsin(x)+x^2 }{ \cos(x) }\] Separate: \[\int\limits \frac{ 2xsin(x) }{ \cos(x) }dx + \int\limits \frac{ x^2 }{ \cos(x) }dx\] which is basically integration by parts for BOTH \[2 \int\limits xtan(x)dx + \int\limits x^2 \sec(x)dx\]

OpenStudy (anonymous):

It supposed to be \[\cos ^{2}x\] for the denominator

OpenStudy (abb0t):

I understand, but one cos(x) get's cancel'd out for the first one. because you have: \[\frac{ 2x \sin(x)\cos(x) }{ \cos^2(x) } = \frac{ 2xsin(x) }{ \cos(x) }\]